OpenStudy (anonymous):

Asia Desai deposited $6,000 in a savings account that pays 5.5% interest compounded daily. How much interest did she earn in 21 days?

OpenStudy (anonymous):

@agent0smith :) ??

OpenStudy (anonymous):

so is it continuously or what?

OpenStudy (anonymous):

wait so compounded daily means once a day?

OpenStudy (anonymous):

I think it is 22909...

OpenStudy (anonymous):

so im guessing it's compounded once a day so 6000x.005=330+6000=6330 and then continue that pattern 21 times... but don't worry I'll do the math.

OpenStudy (anonymous):

there is a shorter way to this but i forgot how to do it...

OpenStudy (anonymous):

It's okay. I got it. Thank you. @helplesspenguin

OpenStudy (anonymous):

Thank you. @VansSkater

OpenStudy (anonymous):

oh how did you do it? did you use sigma notation?

OpenStudy (agent0smith):

\[\Large A= P \left( 1+\frac{ r }{ n } \right)^{nt}\] n in this case is 365 since it's compounded daily. t would be 21/365 since it's only 21 days of a year. r is the interest rate 0.055. P is $6000